Live GMAT classes with Target Test Prep start Nov 21 and Nov 22. First class free.

See classes
Target Test Prep · GMAT

Choose how you want to prepare

Learn live with an expert or move at your own pace. Every option includes the complete TTP study system.

Newsweek Readers’ Choice Awards 2026 winner: Target Test Prep, #1 Test Prep

Target Test Prep was voted #1 Best Test Prep in the 2026 Newsweek Readers’ Choice™ Awards.

★★★★★5.0563 reviews
GMATLiveTeachStarts Nov 22
Josh Braslow, Target Test Prep GMAT expert
LIVE ONLINE CLASSES

Get Ready for GMAT Test Day Faster with Live Online Classes

with Josh Braslow, 20+ years in test prep

Nov 22 · Josh Braslow
Sun · 9:30 AM to 12:30 PM ET
Included
40 hours of live online classes + 6 months of TTP OnDemand
  • Attend the first class for free
  • Every class is recorded, so you never fall behind
View classes & enroll
Limited seats availableTarget Test Prep
EALiveTeachOnDemandStart anytime
EXECUTIVE ASSESSMENT

Target Test Prep EA OnDemand

Self-paced EA prep. Study on your schedule.

EXECUTIVE ASSESSMENT

with

165+ EA score guarantee
$05-day trial no automatic billing
Schedule
Included
40 hours of live online classes plus six months of access to the complete TTP EA OnDemand course.
  • 165+ EA Score Guarantee
  • 4,100+ Quant, Verbal, and Integrated Reasoning practice questions
  • 400+ hours of in-depth video lessons
  • 3,000+ step-by-step video solutions
View EA class & enroll Start free 5-day trial
Limited cohort · enrollment openTrial includes full course accessTarget Test Prep
GMATOnDemandStart anytime
SELF-PACED MASTERCLASS

Target Test Prep GMAT OnDemand

Complete access from day one. Study on your schedule.

715+ score guarantee
$0to start then $127/mo
  • Personalized study plan and analytics
  • Thousands of lessons and practice questions

probability of picking a coin other than a nickel twice

Expert replies
by alivapriyada » Tue Sep 14, 2010 1:21 pm
In United States currency, a nickel is worth 5 cents, a penny is worth 1 cent, and a dime is worth 10 cents. 100 cents equals one dollar. If a hand purse contains 6 nickels, 5 pennies and 4 dimes, what is the probability of picking a coin other than a nickel twice in a row if the first coin picked is not put back?

8/25

12/35

13/35

9/25

17/25
Join the discussion
Source: — Problem Solving |

by Brian@VeritasPrep » Tue Sep 14, 2010 2:00 pm
Good problem! I love GMAT probability strategy, and quite a bit of it comes into play here. A few questions you should ask yourself include:

1) Does the probability of the first event change the probability for the second event?

Here, it does. We start with 15 coins and we'll draw one out and not replace it, so the second event only has 14 coins. (More on that in a second)

2) Do you need a particular sequence to happen, or will multiple sequences accomplish your goal?

Here there is only one sequence that we need: Not Nickel, then Not Nickel.

Sometimes, however, there are multiple sequences that would get you what you need. If this question were "what is the probability of selecting exactly one nickel, then two sequences would work: Nickel --> Not Nickel; or Not Nickel -->Nickel

In this case, we need to multiply our one sequence:

Not Nickel, Not Nickel

The first has a probability of 9 non-nickels / 15 total coins
The second has a probability of 8 non-nickels / 14 total coins (you'll assume that you got what you wanted on the first draw, as we're multiplying the "new probability" by the 9/15 chance that we get what we wanted the first time)

Then multiply:

9/15 * 8/14 is:
3/5 * 4/7 = 12/35
Brian Galvin
GMAT Instructor
Chief Academic Officer
Veritas Prep

Looking for GMAT practice questions? Try out the Veritas Prep Question Bank. Learn More.
Join the discussion

by goyalsau » Tue Sep 14, 2010 2:24 pm
Brian@VeritasPrep wrote:Good problem! I love GMAT probability strategy, and quite a bit of it comes into play here. A few questions you should ask yourself include:

1) Does the probability of the first event change the probability for the second event?

Here, it does. We start with 15 coins and we'll draw one out and not replace it, so the second event only has 14 coins. (More on that in a second)

2) Do you need a particular sequence to happen, or will multiple sequences accomplish your goal?

Here there is only one sequence that we need: Not Nickel, then Not Nickel.

Sometimes, however, there are multiple sequences that would get you what you need. If this question were "what is the probability of selecting exactly one nickel, then two sequences would work: Nickel --> Not Nickel; or Not Nickel -->Nickel

In this case, we need to multiply our one sequence:

Not Nickel, Not Nickel

The first has a probability of 9 non-nickels / 15 total coins
The second has a probability of 8 non-nickels / 14 total coins (you'll assume that you got what you wanted on the first draw, as we're multiplying the "new probability" by the 9/15 chance that we get what we wanted the first time)

Then multiply:

9/15 * 8/14 is:
3/5 * 4/7 = 12/35
Thanks Brian,
Nice work
Join the discussion